# Scalable Two-Stage Stochastic Optimal Power Flow via Separable Approximation

### v1

@@ IV-A Simulation Set-Up
We use Pyomo [13] to model and Gurobi [23] to solve the optimization problems [^1] All the codes implemented to develop SPAR-OPF will be made publicly available once the paper is accepted.. The EF and PH algorithms are implemented using PySP package in Pyomo. All simulations are carried out on a workstation with 32 GB RAM and an Intel Core i9-10900X CPU @ 3.70GHz.
For simulations, we consider $\overline{N}^{DG}=10$, $\underline{P}^{DG}=33\text{kW}$, $\overline{P}^{DG}=333\text{kW}$, $u^{DG}=2\text{kW}$, and $c_{pu}^{DG}=\$1.01$ [24]. A total budget $b$ of \$1500k is assumed. Note that all these parameters can be easily relaxed based on the user’s preference and available resources. All results are generated by averaging the 25 independent random seeds. Unless otherwise mentioned, SPAR-OPF utilized the epigraph $\textbf{E}_{1}$ formulation.
@@ IV-B Illustration of SPAR-OPF
In this section, we illustrate the performance of SPAR-OPF on the IEEE 123 bus system with 96 and 1200 scenarios. The total load of this system is 3431.3 kW and 1931.9 kvar. A system base of 1 MVA and 4.16 kV is considered.
@@ IV-B1 Case I: With 96 scenarios
Figure 3: Illustration of coordinate functions learning. The $x$-axis shows the decision points and the $y$-axis shows the value of the coordinate function.
@@
Table I presents the results from the SPAR-OPF, EF, and PH where columns “Obj” and “Time” indicate the objective value and time taken to solve the problem, respectively. It is important to mention that all algorithms were implemented on same computing environment. For $\textbf{D}_{2}^{\textbf{v}}$, it is observed that the SPAR-OPF objective value is very close to the true objective value (EF column) with a gap of 0.09 (0.44%). The last row shows the solutions for $\textbf{D}_{2}^{\textbf{p}}$ where the solution obtained from the SPAR-OPF is very close to that of the optimal value from EF with a gap of 0.6 kW (0.93%). Note that the proposed framework gives a better solution than PH for $\textbf{D}_{2}^{\textbf{v}}$, whereas later performs better for $\textbf{D}_{2}^{\textbf{p}}$ for the considered problem; nevertheless, both solutions are close to the respective optimal values.
@@ IV-B2 Case II: with 1200 scenarios
In this section, we illustrate statistical bounds and compare solution quality with EF and PH using a significantly larger number of scenarios, and analyze computational time. Table II presents the results for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ considering 1200 scenarios. It is observed that the extensive formulation for this case has 5.28 millions rows and 2.33 millions columns whereas case I has 0.42 millions rows and 0.186 millions columns, which illustrates the increasing complexity of the problem with an increase in the number of scenarios. The column $N_{s}$ in Table II indicates the batch size taken for each random seed while solving the problem using SPAR-OPF. The columns “Obj”, “LB CI”, and “UB CI” refer to objective value, lower bounds confidence interval and upper bound confidence interval for SPAR-OPF respectively. Similarly, column “BG” refers to the bounds gap with $(1-2\alpha)$, where $\alpha=0.05$, confidence level calculated based on the difference between the upper value of the upper bound confidence interval and the lower value of the lower bound confidence interval. The number within the bracket shows the percentage bounds gap to the objective value. Note that the objective value from SPAR-OPF lies within the lower bound and upper bound confidence interval. For smaller batch sizes, the width of the confidence interval and bounds gap are higher as the batch size is insufficient to capture the uncertainty set. The proposed framework obtains a solution closer, and better than the PH solution, to the optimal solution obtained from the EF. The width of UB CI does not change that much as all scenarios are simulated, and less variability is observed.
It is important to mention that a bounds gap of $x\%$ does not necessarily mean that the objective value of SPAR-OPF is $x\%$ away from the optimal solution. It measures the tightness of bounds at the current solution. Take an instance, for case with $N_{s}=200$, the objective value from SPAR-OPF is only 0.05%, far from the optimal value though the BGis 2.1%. It is observed that the proposed framework took 67.8 seconds per seed, whereas EF and PH took 1.86 hours and 2.89 hours to solve the problem, respectively.
@@ IV-C Scalability assessment with system size
In this section, we test the framework’s scalability considering the 9500-node system, a modified IEEE 8500-node test system, with a total load of 13668 kW + 3765 kvar [27]. All secondary loads were referred to the primary side of the transformer. A system base of 1 MVA and 12.47 kV is assumed. We consider 1200 scenarios. EF would result in 112 million rows and 49.4 million columns, which is significantly challenging for any commercial solver to solve this optimization problem as a single large problem in polynomial time. Thus, EF and PH could not solve this problem as it became intractable using the considered workstation. Table IV and Table V present the results for the problem $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ respectively from SPAR-OPF. The obtained solutions are well within the tight bounds where the bounds gap is less than 3% for the batch size $N_{s}\geq 150$. SPAR-OPF took 5.25 minutes and 5.5 minutes per seed to solve $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ respectively. This illustrates the scalability of proposed framework, making it suitable even for real-time applications under limited computing resources.
@@ IV-D Computational comparison between $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$
In this section, the computational performance between two formulations obtained using $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$ is made. Table VI shows the computational performance of these two formulations for different test cases and objectives. The values corresponding to columns $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ represent the time in seconds per iteration while solving the corresponding problems. The number of rows and columns in the optimization problem for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ is slightly higher due to additional constraint (42) and number for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ is reported. Leveraging the convexity of the value function and using only a few additional continuous variables, $\textbf{E}_{\textbf{1}}$ is observed to perform better (4x less time) than $\textbf{L}_{\textbf{1}}$. Even for the larger system size, $\textbf{E}_{\textbf{1}}$ took only 5 seconds per iteration to solve the problem. Based on all simulations, 80 iterations were required on average for the convergence. This confirms that the proposed framework applies to realistic large-scale power systems.
@@ V Conclusion and Future Work
We developed the SPAR-OPF framework to solve OPF-based two-stage stochastic optimization problems in power systems. To the best of our knowledge, this is the first to solve OPF-based two-stage stochastic optimization problems using separable projective approximations of value function. We provided two formulations for the learned functions and a statistical validation procedure. The proposed framework is illustrated on two standard three-phase unbalanced distribution system test cases considering DG siting and sizing problem. Results illustrated that SPAR-OPF provides near-optimal solutions with significantly lower solve times as compared to state-of-the-art methods. The scalability of the proposed framework is validated with a larger system size and a higher number of scenarios, making it suitable even for resource-constrained environments. As a part of future work, we will explore heuristics such as MIP relaxation, exploration vs. exploitation, and successive MIP gap reduction to further reduce the solution time without affecting the final solution. The proposed framework will further be demonstrated on classical stochastic problems in power systems such as unit commitment and dynamic economic dispatch.

### v2

@@ IV-A Simulation Set-Up
We use Pyomo [13] to model SPAR-OPF, PySP package in Pyomo to model EF and PH, and Gurobi to solve all models [25]. All the codes and data implemented to develop SPAR-OPF, EF and PH will be made publicly available once the paper is accepted. All simulations are carried out on a workstation with 32 GB RAM and an Intel Core i9-10900X CPU @ 3.70GHz.
For simulations, we consider $\overline{N}^{DG}=10$, $\underline{P}^{DG}=33\text{kW}$, $\overline{P}^{DG}=333\text{kW}$, $u^{DG}=2\text{kW}$, and $c_{pu}^{DG}=\mathdollar 1.01$ [26]. A total budget $b$ of \$1500k is assumed. Note that all these parameters can be easily relaxed based on the user’s preference and available resources. All results are generated by averaging the 25 independent random seeds. Unless otherwise mentioned, SPAR-OPF utilized the epigraph $\textbf{E}_{1}$ formulation.
@@ IV-B Illustration of SPAR-OPF
In this section, we illustrate the performance of SPAR-OPF on the IEEE 123 bus system with 96 and 1200 scenarios. The total demand is 3431.3 kW and 1931.9 kvar. We consider a system base of 1 MVA and voltage base of 4.16 kV.
@@ IV-B1 Case I: With 96 scenarios
Figure 2: Illustration of coordinate functions learning.
@@
Table I presents the results from SPAR-OPF, EF, and PH where columns “Obj” and “Time” indicate the objective value and time taken to solve the problem, respectively. It is important to mention that all algorithms were implemented on same computing environment. For $\textbf{D}_{2}^{\textbf{v}}$, it is observed that SPAR-OPF objective value is very close to the true objective value (EF column) with a gap of 0.09 (0.44%). The last row shows solutions for $\textbf{D}_{2}^{\textbf{p}}$ where the solution obtained from SPAR-OPF is very close to that of optimal value from EF with a gap of 0.6 kW (0.93%). Note that the proposed framework gives a better solution than PH for $\textbf{D}_{2}^{\textbf{v}}$, whereas later performs better for $\textbf{D}_{2}^{\textbf{p}}$ for the considered problem; nevertheless, both solutions are close to the respective optimal values.
@@ IV-B2 Case II: with 1200 scenarios
In this section, we illustrate statistical bounds and compare solution quality with EF and PH using a significantly larger number of scenarios, and analyze computational time. Table II presents the results for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ considering 1200 scenarios. It is observed that EF for this case has 5.28 millions rows and 2.33 millions columns whereas case I has 0.42 millions rows and 0.186 millions columns, which illustrates the increasing complexity of the problem with an increase in the number of scenarios. The column $N_{s}$ in Table II indicates the batch size taken for each random seed while solving the problem using SPAR-OPF. The columns “Obj”, “LB CI”, and “UB CI” refer to objective value, lower bounds confidence interval and upper bound confidence interval for SPAR-OPF respectively. Similarly, column “BG” refers to the bounds gap with $(1-2\alpha)$, where $\alpha=0.05$, confidence level calculated based on the difference between the upper value of the upper bound confidence interval and the lower value of the lower bound confidence interval. The number within the bracket shows the percentage bounds gap to the objective value. Note that the objective value from SPAR-OPF lies within the lower bound and upper bound confidence interval. For smaller batch sizes, the width of the confidence interval and bounds gap are higher as the batch size is insufficient to capture the uncertainty set. The proposed framework obtains a solution closer, and better than the PH solution, to the optimal solution obtained from the EF. The width of UB CI does not change that much as all scenarios are simulated, and less variability is observed.
It is important to mention that a bounds gap of $x\%$ does not necessarily mean that the objective value of SPAR-OPF is $x\%$ away from the optimal solution. It measures the tightness of bounds at the current solution. Take an instance, for case with $N_{s}=200$, the objective value from SPAR-OPF is only 0.05%, far from the optimal value though the BG% is 2.1%. It is observed that the proposed framework took 67.8 seconds per seed, whereas EF and PH took 1.86 hours and 2.89 hours to solve the problem, respectively.
@@ IV-C Scalability assessment with system size
We test the framework’s scalability on the 9500-node system, a modified IEEE 8500-node test system, with a total load of 13,668 kW + 3,765 kvar [29], using a system base of 1 MVA and a voltage base of 12.47 kV. 1200 scenarios are considered. EF would result in 112 million rows and 49.4 million columns, which is significantly challenging for any commercial solver to solve this optimization problem as a single large problem in polynomial time. Thus, EF and PH could not solve this problem as it became intractable using the considered workstation. Table IV and Table V present the results for the problem $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ respectively from SPAR-OPF. The obtained solutions are well within the tight bounds where the bounds gap is less than 3% for the batch size $N_{s}\geq 150$. SPAR-OPF took 5.25 minutes and 5.5 minutes per seed to solve $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ respectively. This illustrates the scalability of proposed framework, making it suitable even for real-time applications under limited computing resources.
@@ IV-D Computational comparison between $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$
This section compares the computational performance of two formulations, $\textbf{L}_{\textbf{1}}$ and $\textbf{E}_{\textbf{1}}$. Table VI reports results for different test cases and objectives, where columns $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ and $\textbf{D}_{\textbf{2}}^{\textbf{p}}$ denote per-iteration solve times. The number of rows and columns for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ is slightly higher due to additional constraint (43) and number for $\textbf{D}_{\textbf{2}}^{\textbf{v}}$ is reported. Leveraging the convexity of value function and using only a few additional continuous variables, $\textbf{E}_{\textbf{1}}$ is observed to perform better (4x less time) than $\textbf{L}_{\textbf{1}}$. Even for larger system size, $\textbf{E}_{\textbf{1}}$ required only 5 seconds per iteration. Based on all simulations, 80 iterations were required on average for convergence. This confirms that the SPAR-OPF applies to realistic large-scale power systems.
@@ V Conclusion and Future Work
We developed the SPAR-OPF framework to solve OPF-based two-stage stochastic optimization problems in power systems. To the best of our knowledge, this is the first application of separable projective approximations of the value function to such problems. We provided two formulations for learned functions and a statistical validation procedure. The framework is demonstrated on two standard three-phase unbalanced distribution system test cases for DG siting and sizing problem, showing near-optimal solutions with significantly lower solve times compared to state-of-the-art methods. The scalability of the framework is validated on a larger system with a higher number of scenarios, making it suitable even for resource-constrained environments. As future work, we will explore heuristics such as MIP relaxation, exploration vs. exploitation, and successive MIP gap reduction to further reduce solution time and extend the framework to classical stochastic problems such as unit commitment and dynamic economic dispatch.